# Enhancing Inertial Hand based HAR through Joint Representation of Language, Pose and Synthetic IMUs

### v1

@@ 4.3. Discussion and Limitations
To investigate the impact of different components on performance, we conduct experiments with varied settings.
Effect of Pretraining. Pretraining with simulated data in fact helps in creating better clustering of different latent representations. As we can see in Figure 3, the combination of pretraining followed by fine-tuning yields the highest discriminability of latent representations in IMU data. While some clustering is observable in Figures 3 (2) and (4), these representations appear less structured compared to Figure 3 (5). In contrast, Figure 3 (1) illustrates scattered clusters, suggesting that raw IMU data without pretraining and fine-tuning lacks clear patterns in latent representation. Even if we are pretraining on a rich dataset that is not of the target it still improves the overall results. Also, our method, not frozen, consistently improves on our baseline, which is not always true for other simulation methods. Additionally based on the results we can clearly see that single-task pretrainings like Contrastive pretrain and IMU reconstruction are outperformed by multi-task pretraing used in $\text{Multi}^{3}$et regardless of the type of pretraining or downstream dataset. Hence proving our hypothesis multi-task pretraining is superior.
Effect of Freezing Encoder Weights Based on the results presented in the tables, the ”not frozen” approach outperforms the ”frozen” approach on both datasets. This indicates that optimizing the encoders pretrained from a different data modality (i.e., video) is necessary for IMU-based HAR tasks. As the pretraining dataset share non standard activity types and does not really contains both the specific activities present in the target dataset as well as the calibrated imu signals that falls in same range as target IMU signal it is expected to have poorer result when the encoder is not learning anything from the target dataset. Nevertheless this method help us in improving the overall HAR when we don’t freeze it as it has acquired some knowledge during pretraing process and when that mixed with knowledge acquired during downstream training making it better than other methods.
@@
Figure 4. F1-score on different amount of real IMU used for downstream task (left wrist at top and both wrists at bottom).
We conduct experiments on OpenPack dataset to investigate how the size of the training set influences the macro F1-score of our approach. As depicted in Figure 4, as the amount of real IMU data increases, the macro F1-score of the proposed approach also increases. However, while the result of utilizing synthetic IMU data directly with only 25% real IMU data is the highest, the synthetic IMU data is generated from the video of the target dataset. This is in contrast to the proposed method, where the activities in the video are completely different from the target dataset.
Limitations. As shown in the tables, different datasets for pretraining have an impact on the downstream HAR performance, highlighting the importance of selecting appropriate datasets for pretraining. Furthermore, it is observed that the performance on relatively simple activities does not show a significant increase. This lack of enhancement can likely be attributed to the fact that the pretraining datasets predominantly consist of non-standard activity types, which may not effectively generalize to more straightforward activities. This discrepancy suggests that the pretraining process might not be adequately capturing the nuances required for these simpler tasks. Additionally, as illustrated in Figure 1, the simulated IMU data exhibits notable imperfections, whose severity correlates with the accuracy of the estimated pose. This dependence underscores the critical role of pose estimation quality in shaping the fidelity of simulated IMU data. Consequently, the downstream performance of HAR models may be affected by the fidelity of the simulated IMU data, necessitating careful consideration and refinement of pose estimation techniques in conjunction with HAR model development.
Our model performs comparatively worse than other approaches for single wrist-based IMU detection. This discrepancy may be due to the data interpretation during pretraining, which utilized both wrists, whereas downstream training involved only one wrist. To address this issue, a more comprehensive study and additional experiments are needed.
@@ 5. Conclusion
This paper presents $\text{Multi}^{3}$et, an extensive framework that leverages multi-modal contrastive learning to enhance IMU-based HAR performance using video data (sign language), with activity classes distinct from the target datasets. The IMU simulation approach is utilized to effectively extract fine-grained activity features from the video, and the learned features enhance downstream HAR tasks, particularly in recognizing complex activities with limited data availability. Experimentation conducted for fine-tuning across various settings demonstrates that the ”no frozen” setting consistently outperforms the others, exhibiting an 8 % improvement compared to baseline methods. In the future, we plan to investigate the impact of pretraining data selection, the performance on single wrist IMUs, and the challenge of accurately estimating poses for high-fidelity IMU simulation.

### v2

@@ 4.3. Discussion and Limitations
Effect of Pretraining. Pretraining with simulated data in fact helps in creating better clustering of different latent representations. As we can see in Figure 3, the combination of pretraining followed by fine-tuning yields the highest discriminability of latent representations in IMU data. While some clustering is observable in Figures 3 (2) and (4), these representations appear less structured compared to Figure 3 (5). In contrast, Figure 3 (1) illustrates scattered clusters, suggesting that raw IMU data without pretraining and fine-tuning lacks clear patterns in latent representation. Even if we are pretraining on a rich dataset that is not of the target it still improves the overall results. Also, our method, not frozen, consistently improves on our baseline, which is not always true for other simulation methods. Additionally based on the results we can clearly see that single-task pretrainings like Contrastive pretrain and IMU reconstruction are outperformed by multi-task pretraining used in $\text{Multi}^{3}$et regardless of the type of pretraining or downstream dataset. Hence demonstrating how multi-task pretraining is superior to the single-task approach.
Effect of Freezing Encoder Weights Based on the results presented in the tables, the ”not frozen” approach outperforms the ”frozen” approach on both datasets. This indicates that optimizing the encoders pretrained from a different data modality (i.e., video) is beneficial for IMU-based HAR tasks. This is to be expected as the pretraining datasets share non standard activity types and do not really contains both the specific activities present in the target dataset. Given differences regarding IMU calibration and signal ranges indeed we expected poorer results when the encoder is not learning anything from the target dataset. Nevertheless $\text{Multi}^{3}$et can improve the overall HAR when we allow encoders that acquired some knowledge during the pretraining process to be later fine-tuned with knowledge related to the downstream task, with this approach consistently outperforming other methods.
@@
Figure 4. Macro F1-score on different amount of IMU used for downstream task (left wrist: top,both wrists: bottom) using Baseline, DCL (only real data, real+virtual data IMUTube) and pretrained $\text{Multi}^{3}$et .
We conduct experiments on the OpenPack dataset to investigate how the size of the training set influences the macro F1-score of our approach. As depicted in Figure 4, our method is label efficient, with similar or better performance than the full training set with 25% of it. As the amount of real IMU data increases, the macro F1-score of the proposed approach also increases, as opposed to virtual data generation approaches, that are close or worse than their baselines at higher data rates.
Limitations. As shown in the tables, different datasets for pretraining have an impact on the downstream HAR performance, highlighting the importance of selecting appropriate datasets for pretraining. Furthermore, it is observed that the performance on relatively simple activities does not show a significant increase. This lack of enhancement can likely be attributed to the fact that the pretraining datasets predominantly consist of non-standard activity types, which may not effectively generalize to more straightforward activities. This discrepancy suggests that the pretraining process might not be adequately capturing the nuances required for these simpler tasks. Additionally, as illustrated in Figure 1, the simulated IMU data exhibits notable imperfections, whose severity correlates with the accuracy of the estimated pose. This dependence underscores the critical role of pose estimation quality in shaping the fidelity of simulated IMU data. Consequently, the downstream performance of HAR models may be affected by the fidelity of the simulated IMU data, necessitating careful consideration and refinement of pose estimation techniques in conjunction with HAR model development. We also need further study on the effect of the Base model. Our performance may also be currently limited by the baseline architecture, as it alone does not outperform DCL in the OpenPack dataset.
@@ 5. Conclusion
This paper presents $\text{Multi}^{3}$et, an extensive framework that leverages multi-modal contrastive learning to enhance IMU-based HAR performance using video data (sign language), with activity classes distinct from the target datasets. The IMU simulation approach is utilized to effectively extract fine-grained activity features from the video, and the learned features enhance downstream HAR tasks, particularly in recognizing complex activities with limited data availability. Experimentation conducted for fine-tuning demonstrates that the ”not frozen” setting consistently outperforms the others, exhibiting from 5% to 18% improvement compared to baseline methods. In the future, we plan to investigate the impact of pretraining data selection, the challenge of accurately estimating poses for high-fidelity IMU simulation as well as comparing our method to other multi-modal pretraining approaches such as MESEN(Xu et al., 2024).
Acknowledgements.
This paper was supported by the BMBF (German Federal Ministry of Education and Research) in the project VidGenSense (01IW21003), the Carl Zeiss Stiftung under the Sustainable Embedded AI project (P2021-02-009) and China NSFC Grant U2001207, Guangdong Provincial Key Lab of Integrated Communication (No.2023B1212010007), the Project of DEGP (No.2023KCXTD042).